Cindy’s tennis club raises its price of an hour of tennis from $5 to $10, other things remaining 

the same.

6- List the combinations of hours spent playing golf and tennis that Cindy can now afford 

and her marginal utility per dollar from golf and from tennis.

7- How many hours does Cindy now spend playing golf and how many hours does she 

spend playing tennis?

8- Use the information in Problem 6 to draw Cindy’s demand curve for tennis. Over the 

price range of $5 to $10 an hour of tennis, is Cindy’s demand for tennis elastic or 

inelastic?

9- Explain how Cindy’s demand for golf changed when the price of an hour of tennis 

increased from $5 to $10 in problem 6


1
Expert's answer
2021-06-24T12:04:30-0400

6:

Given income received by Cindy is $ 70. Per hour of tennis. With a value of 10, the maximum number he can play tennis is 7 (and no golf). Likewise, she can play 7 hours of golf (and no tennis) with the given income.

Bundle is affordable when it can satisfy the budget constraint

"M=xpx+ypy\\m=income\\x=good\\y=good"

"px" =price of good x and "py" =price of good y

so substitute the values of the connection in the disposable bundle

price of two goods and the given income to check whether disposable bundle is affordable

"70=xpx+xpx\\70=T\times 10+G\times 10\\7=T+G"


Therefore, the budget limit suggests that all the bundles of hours spent on tennis and golf, which satisfy, are affordable. Affordable bundles are: 0 hours of tennis and 7 hours of golf, 1 hour of tennis and 6 hours of golf, 2 hours of tennis and 5 hours of golf, 3 hours of tennis and 4 hours of golf, 4 hours of tennis and 3 hours of golf, 5 hours of tennis and 2 hours of golf, 6 hours of tennis and 1 hour of golf Hours and 7 hours of tennis and 0 hours of golf.


7:

Hourly Rate Gold $10

Hourly Rate Tennis  $10




To examine the table note that the marginal utility per dollar spent in the last hour of tennis is the same as the marginal utility per dollar spent in the last hour of golf when Cindy chooses last hour tennis and hours golf. This combination is on the budget line ($ 10 × 3 + 10 × 4 = $ 70).


8:





9:

Cross elasticity"=\frac{∆Q_g}{∆p_t}×\frac{p_t}{Q_g}"

Change in quantity of golf and change in price of tennis

"∆Qs=4-5=1\\∆p.10-5=5\\Q_g=5,P_t=5\\cross\space elasticity=\frac{-1}{5}×\frac{5}{5}\\=-0.2"

Golf and tennis are compliments.demand for golf falls if price of tennis is increased.